OpenStudy (anonymous):

Write the equation of the line given by the following graph in standard form. graph of a line in the coordinate plane with a slope of negative two and points at (–1, 3) and (1, –1) 2x – y = –1 2x + y = –1 2x + y = 1 2x – y = 1

OpenStudy (aravindg):

whats the difficulty u have here ?

OpenStudy (anonymous):

Well, I think the answer is 2x+y=1 But, I'm not sure. :/

OpenStudy (aravindg):

using slope point form with slope -2 and point(-1,3) y-3 =-2(x+1) y-3=-2x-2 2x+y=1 yep u are right !!!

OpenStudy (anonymous):

Thank you :)